# Deep Support Vectors

### v1

@@ Abstract
While the success of deep learning is commonly attributed to its theoretical equivalence with Support Vector Machines (SVM), the practical implications of this relationship have not been thoroughly explored. This paper pioneers an exploration in this domain, specifically focusing on the identification of Deep Support Vectors (DSVs) within deep learning models. We introduce the concept of DeepKKT conditions, an adaptation of the traditional Karush-Kuhn-Tucker (KKT) conditions tailored for deep learning. Through empirical investigations, we illustrate that DSVs exhibit similarities to support vectors in SVM, offering a tangible method to interpret the decision-making criteria of models. Additionally, our findings demonstrate that models can be effectively reconstructed using DSVs, resembling the process in SVM. The code will be available.
Machine Learning, ICML
@@ 1 Introduction
Deep learning has led revolutionary advancements in modern technology, catalyzing paradigm shifts across numerous fields. However, despite these successes, a fundamental understanding of how deep learning models excel in pattern generalization and the criteria they use for decision making remains uncertain. This research endeavors to demystify the black box of deep learning, offering deeper insights into how these models process and interpret data.
In explaining the performance of deep learning, one of the most useful attempts is associating it with support vector machines (SVM) (Cortes & Vapnik, 1995; Hearst et al., 1998) because SVM theoretically provides a well-tight bound on its performance. Consequently, support vectors, which encode the decision boundary, offer valuable insights into how the model classifies based on certain criteria. Moreover, the number of support vectors is much smaller than the size of the trained dataset, making it parameter-efficient. This idea aligns well with deep learning models that converge to 100% training accuracy in overparameterized settings (Zhang et al., 2021). It has been proven that feedforward networks, including convolutional networks, can be equivalent to SVM (Lyu & Li, 2020; Soudry et al., 2018; Chen et al., 2021) and there is an active effort to bring the SVM architecture itself into the realm of deep learning models (Wang et al., 2019; Tang, 2013).
Hence, it can be logically inferred that a deep learning model, akin to an SVM, should demonstrate comparable fundamental traits. Specifically, the model is expected to have a corresponding set of support vectors that define its decision boundary. However, attempts to extract support vectors from generally trained deep learning models have not yet been made.
In this paper, we focus on the crucial issue of the practical relationship between support vectors and deep learning. Based on the theoretical equivalence (Soudry et al., 2018), our research represents the first attempt to connect traditional machine learning, epitomized by SVM, with typical deep learning models. In doing so, we define ‘Deep Support Vectors’ (DSVs) as the samples constituting decision boundary in the input space and their desirable characteristics analogous to their counterparts in linear SVMs. Reinterpreting the traditional SVM’s KKT (Karush-Kuhn-Tucker) condition (Bertsekas, 1999) for deep networks, we introduce ’DeepKKT’, a revised KKT condition tailored for deep learning, along with the concept of Deep Support Vectors. DeepKKT smoothly blends the principles of classical machine learning with the deep learning realm, making it compatible with the deep learning feature of handling high-dimensional data that reside on a low-dimensional manifold.
Figure 1: Comparison of synthesized images (first row) created using the DeepKKT condition initiated from noise, and selected images (second row) from the CIFAR-10 training dataset. The selected images were chosen based on $\lambda$ values, i.e., the ten images have the highest $\lambda$ in each class. Both synthesized and selected images demonstrate similarity at the pixel level sharing common features.
In our experiments with DSVs, we utilize DeepKKT in two ways. Firstly, we select DSVs from the training dataset using DeepKKT as a scoring function measuring the closeness of a sample to the decision boundary. Secondly, we synthesize DSVs from pure noise, employing DeepKKT as a loss function in model inversion settings (See Fig. 1).
In this study, we performed experiments to elucidate the attributes of synthesized DSVs. Our findings reveal that DSVs consistently meet all conditions analogous to support vectors in SVMs. Additionally, we demonstrated that DSVs contain significantly denser information than other data samples in the training dataset. Furthermore, we demonstrated that modifying the predictions of the model by editing the information gathered from DSVs is possible. This illustrates that DSVs can play a role in elucidating the decision criteria of the model. Also, we demonstrated that DSVs can actually serve as the same role as support vectors, meaning than we can retrain the model up to a level using only a single image per category. i.e., we used DSVs in dataset distillation (Cazenavette et al., 2022; Wang et al., 2018; Liu et al., 2022). Our performance is comparable to others for preexisting benchmarks . Also, our setting is far more reasonable than others because they demand substantial computational resources as they normally use gradient matching between the distilled dataset and the entire training dataset. This requires heavy computational load, which is paradoxical as their goal is to reduce computation burden while maintaining privacy.
In contrast to existing studies that solely provide input-dependent explanations for decision-making, DSVs are the first method capable of elucidating the decision boundary’s criterion solely through the trained model, thereby unraveling the black-box nature of deep learning.
@@ 2.1 SVM in Deep Learning
Numerous studies have endeavored to establish a connection between deep learning and conventional SVMs. Recently, Soudry et al. (2018) demonstrated that, in an overparameterized setting with linearly separable data, the learning dynamics of a model possessing a logistic tail are equivalent to those of a support vector machine, with the model’s normalized weights converging towards a finite value. Following this, Lyu & Li (2020) extended this equivalence to feedforward networks. Other studies (Wang et al., 2019; Tang, 2013) directly integrated SVM into their architecture to capitalize on its beneficial properties. Despite these advancements, there is still a lack of research that directly explores support vectors through this theoretical lens of equivalence. In this study, we address this theoretical equivalence of support vectors by introducing the DeepKKT condition, a KKT condition tailored for deep learning, thus applying this equivalence in a practical context.
@@ 2.2 Model Inversion Through the Lens of Maximum Margian
There is a line of research utilizing the stationarity condition, a part of the KKT condition, for model inversion. Haim et al. (2022) firstly exploited the KKT condition for model generation, adopting SVM-like architectures. They conducted experiments with binary classification, a 2-layer MLP, and full-batch gradient descent. Buzaglo et al. (2023) extended these experiments to multi-label classification by adapting their existing architecture to a multi-class SVM structure. To ensure the generated samples lie on the data manifold, they were initialized with the dataset’s mean, implying the adoption of some prior knowledge of the data. Similarly, Yu & Wang (2023) generated images through the stationarity condition, also adopting the mean-initialization and conducting experiments on the MNIST (Deng, 2012) and CelebA (Liu et al., 2015) datasets.
Contrary to previous studies, our research extensively employs the KKT conditions, not just focusing on the stationarity condition. While previous works adopted SVM-like structures to directly utilize the stationarity condition, we derived DSVs from universal deep learning settings including SGD (Stochastic Gradient Descent), mini-batch processing, multi-class classification, and data augmentation, indicating that DSVs can be obtained in any trained model settings, not just within SVM-like architectures. Furthermore, their limited number aligns with the core principle of support vectors. These distinctions underscore the unique contributions of our work. This was possible because we reinterpreted the KKT condition used in machine learning to fit in deep learning settings.
@@ 2.3 Dataset Distillation
Dataset distillation (Cazenavette et al., 2022; Wang et al., 2018; Liu et al., 2022) fundamentally aims to reduce the size of the original dataset while maintaining model performance. The achievement also involves addressing privacy concerns and alleviating communication issues between server nodes. Although this approach does not explicitly clarify the link with SVM, distilled datasets share key features with support vectors in that they enable model reconstruction with minimal data. The problem is typically addressed under the following conditions: 1) Access to the entire dataset for gradient matching, 2) Possession of snapshots from all stages of the model’s training phase, and 3) Availability of multiple models with different initializations. These conditions present greater complexity and constraints relative to our DSV extraction, posing practical implementation challenges in real-world situations. In contrast, our algorithm does not require all three of these conditions and operates based just on a single trained model.
@@ 3.1 Notation
In SVM formulation, $\tilde{w}(:=[w;b])$ represents the concatenated weight vector $w$ and bias $b$. Each data instance, expanded to include the bias term, is denoted by $\tilde{x}_{i}(:=[x_{i};1])$, while the corresponding binary label is represented by $y_{i}\in\{\pm 1\}$. The Lagrange multipliers are denoted by $\alpha_{i}$’s.
Transitioning to the context of deep learning, we denote the parameter vector of a neural network model by $\theta$, which, upon optimization, yields $\theta^{*}$ as the set of learned weights. The mapping function $\Phi(x_{i};\theta)$ represents the transformation of input data into a $C$-dimensional logit in a $C$-class classification problem in a manner dictated by the parameters $\theta$, i.e., $\Phi(x_{i};\theta)=[\Phi_{1}(x_{i};\theta),\cdots,\Phi_{C}(x_{i};\theta)]^{T}\in\mathbb{R}^{C}$. we define score as the logit of a target index. For example, $\Phi_{y_{i}}(x_{i};\theta)$ is the value of the logit for the target class. If the score is the largest among logits, i.e., $\operatorname*{argmax}_{c}\ \Phi_{c}(x_{i};\theta)=y_{i}$, then it correctly classifies the sample. The Lagrange multipliers adapted to the optimization in deep learning are represented by $\lambda_{i}$’s. $(x,y)$ denotes a pair of input and output and $\mathcal{I}$ is the index set.
This section presents the specific conditions that DSVs (Deep Support Vectors) must satisfy and discusses how to achieve that condition within gradient descent framework.
@@ 3.2 Support Vector Machines
The fundamental concept of Support Vector Machines (SVMs) is to find the optimal hyperplane that classifies the given data. This hyperplane is defined by the closest data points to itself known as support vectors, and the distance between the support vectors and the hyperplane is termed the margin. The hyperplane must classify the classes correctly while maximizing the margin. This leads to the following conditions that an SVM must satisfy:
Primal feasibility: $\forall i,\ y_{i}\tilde{w}^{T}\tilde{x}_{i}\geq 1,$ (1)
Dual feasibility: $\forall i,\ \alpha_{i}\geq 0,$
[…]
Stationarity: $\tilde{w}=\sum_{i=1}^{n}\alpha_{i}y_{i}\tilde{x}_{i}.$
The primal and dual conditions ensure that these critical values correctly classify the data while being outside the margin. The complementary slackness condition mandates that support vectors lie on the decision boundary. The stationarity condition makes sure that the derivative of the Lagrangian be zero.
The final condition, stationarity, offers profound insights into Support Vector Machines (SVM). It underscores the fact that support vectors are integral to encoding the decision boundary $\tilde{w}$. Consequently, identifying the decision hyperplane in SVMs is tantamount to pinpointing the corresponding support vectors. This implies that with a trained model at our disposal, we can reconstruct the SVM in two distinct ways:
1. Support Vector Selection: From the trained model, we can extract support vectors that inherently encode the decision hyperplane.
@@ 4 Deep Support Vector
This section presents the specific conditions that DSVs (Deep Support Vectors) must satisfy and discusses how to get an optimization loss with low computation cost that meets these conditions.
Algorithm 1 Support Vector Refinement for Deep Learning Model
  Input: Pretrained model $\Phi(\cdot;\theta)$, loss function $L$, augmentation function set $A$, number of DSV candidate $N$, number of class $C$, hyperparameter $\beta$, $\gamma$, $\alpha_{\lambda}$, $\alpha_{x}$
  Ensure: Freeze model $\Phi(\cdot;\theta)$
  Initialize $N\times C$ number of support vector candidates
  for $c=1$ to $C$ do
     Sample $N$ number of ($x^{c}_{i}$,$\lambda^{c}_{i}$) for label $c$
  end for
  Define $X=\{x^{c}_{i}\mid c\in[C],i\in[N]\}$
  Define $\Lambda=\{\lambda^{c}_{i}\mid c\in[C],i\in[N]\}$
  repeat
     $L_{\text{primal}}(X)=(\sum_{i=1}^{N}\sum_{c=1}^{C}L(\Phi(x^{c}_{i};\theta),c))/NC$
     $L_{\text{stat}}(X)=||\theta+\sum_{i=1}^{N}\sum_{c=1}^{C}\lambda^{c}_{i}\nabla_{\theta}L(\Phi(x^{c}_{i};\theta),c)||^{2}_{2}$
     $L_{\text{kkt}}(X)=L_{\text{primal}}(X)+\beta\times L_{\text{stat}}(X)$
     Sample $f_{A}\in A$
     Define $X^{\prime}=\{f_{A}(x^{c}_{i})\mid x^{c}_{i}\in X\}$
     ${L_{total}(X)}={L_{kkt}(X)}+\gamma\times{L_{kkt}(X^{\prime})}$
     Update $x_{i}^{c}\leftarrow x_{i}^{c}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha_{x}}\nabla_{x_{i}^{c}}L_{total}(X)$,  $\forall\ i\ \text{and}\ c$
     Update $\lambda_{i}^{c}\leftarrow\lambda_{i}^{c}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha_{\lambda}}\nabla_{\lambda_{i}^{c}}L_{total}(X)$,  $\forall\ i\ \text{and}\ c$
     Remove $x^{c}_{i}$ from $X$ if corresponding $\lambda^{c}_{i}<0$
  until $X$ converges
  Return: Set of DSV : $X$
@@ 4.1 DeepKKT
$\begin{split}&\text{Hinge Loss:}\ L_{h}(x_{i},y_{i};\tilde{w})=\frac{1}{n}\sum_{i=1}^{n}\max(0,1-y_{i}(\tilde{w}^{T}\tilde{x}_{i})),\\ &\quad\nabla_{\tilde{w}}L_{h}=\frac{1}{n}\sum_{i=1}^{n}\begin{cases}0,&\text{if }y_{i}(\tilde{w}^{T}\tilde{x}_{i})>1\\ -y_{i}\tilde{x}_{i},&\text{otherwise.}\end{cases}\end{split}$ (2)
With Eq. (LABEL:eq:hingegrad), the stationarity condition in Eq. (1) becomes
$w^{*}:=-\sum_{i=1}^{n}\alpha_{i}\nabla_{w}L(x_{i},y_{i};w^{*}),\quad\text{s.t.}\ \alpha_{i}>0.$ (3)
@@
In this paper, we adapt the KKT conditions presented in Eq. (1) to deep learning. In doing so, the two main hurdles of a deep network different from a linear binary SVM are 1) the nonlinearity of $\Phi(x_{i};\theta)$ taking the role of $\tilde{w}^{T}\tilde{x}_{i}$ and 2) multi-class nature of a deep learning model.
Considering the role of the primary condition is to correctly classify $x_{i}$ into $y_{i}$, we can enforce the score $\Phi_{y_{i}}(x_{i};\theta)$ for the correct class $y_{i}$ to take the maximum value among all the logits with some margin $\epsilon$, i.e.,
@@
Because this condition is discontinuous, we can relax it with a surrogate function $-L(\Phi(x_{i};\theta^{*}),y_{i})$, which is the negative loss function to maximize. Note that if we take the cross-entropy loss for $L$, it becomes
@@
$\theta^{*}=-\sum_{i=1}^{n}\lambda_{i}\nabla_{\theta}L(\Phi(x_{i};\theta^{*}),y_{i}).$ (6)
This condition is also a generalized formulation of Eq. (3), where we substitute the linear model $\tilde{w}^{T}\tilde{x}$ with a nonlinear model $\Phi(x;\theta)$, and the binary classification hinge loss $L_{h}$ with multi-class classification loss $L$. Furthermore, the stationarity condition reflects the dynamics of overparameterized deep learning models. We provide an analogy in Appendix A.
However, these conditions are not enough for deep learning. As mentioned before, we are interested in synthesizing support vectors. Compared to the problems dealt with in the classical SVMs, the input dimensions of deep learning problems are typically much higher. In this case, the data are likely to lie along a low-dimensional latent manifold $\mathcal{M}$ inside the high-dimensional space. To ensure a generated sample be a plausible DSV, it not only should satisfy the generalized KKT condition but also should lie in an appropriate data manifold, i.e., $x\in\mathcal{M}$.
@@
Primal feasibility: $\forall\,i\in\mathcal{I},\quad\operatorname*{argmax}_{c}\;\Phi_{c}(x_{i};\theta^{*})=y_{i}$
@@
4.2 Deep Support Vectors, From Dust to Glory
In Sec, 4.1, we explored the KKT conditions in the context of deep learning and how these conditions can be used to formulate a loss function. It is important to note that our goal is not to construct the model $\Phi$, but rather to generate support vectors of an already-trained model $\Phi$ with its parameter $\theta^{*}$.
Our paper addresses the issue of reconstructing support vectors from trained deep learning models. Essentially, this involves sampling or synthesizing support vectors. By replacing the optimization variable $\theta$ with $x$, we shift our focus to utilizing the DeepKKT conditions for generating or evaluating input $x$ rather than $\theta$. This adjustment necessitates a consideration of the data’s inherent characteristics, specifically its multiclass nature and the tendency of the data to reside on a lower-dimensional manifold compared to its defined dimension.

### v2

@@ Abstract
Deep learning has achieved tremendous success. However, unlike SVMs, which provide direct decision criteria and can be trained with a small dataset, it still has significant weaknesses due to its requirement for massive datasets during training and the black-box characteristics on decision criteria. This paper addresses these issues by identifying support vectors in deep learning models. To this end, we propose the DeepKKT condition, an adaptation of the traditional Karush-Kuhn-Tucker (KKT) condition for deep learning models, and confirm that generated Deep Support Vectors (DSVs) using this condition exhibit properties similar to traditional support vectors. This allows us to apply our method to few-shot dataset distillation problems and alleviate the black-box characteristics of deep learning models. Additionally, we demonstrate that the DeepKKT condition can transform conventional classification models into generative models with high fidelity, particularly as latent generative models using class labels as latent variables. We validate the effectiveness of DSVs using common datasets (ImageNet, CIFAR10 and CIFAR100) on the general architectures (ResNet and ConvNet), proving their practical applicability. (See Fig. 1)
Figure 1: Generated images using a model trained with the ImageNet, CIFAR10, and CIFAR100 datasets, respectively. Each image was generated without referencing the original training data.
@@ 1 Introduction
Although deep learning has gained enormous success, it requires huge amounts of data for training, and its black-box characteristics regarding decision criteria result in a lack of reliability. For example, CLIP [24] needs 400 million image pairs for training and Stable Diffusion XL (SDXL) [25] requires 5 billion images. This implies only a small number of groups can train foundation models from scratch. Also, the black-box nature makes it hard to anticipate the model’s performance in different environments. For example, suppose we are to classify pictures of deer and most training deer images contain antlers. For the test images taken in winter, the performance will be worse as deer shed their antlers in winter. As modern deep learning models do not provide any decision criterion i.e., black box, we cannot determine whether the domain of the model has shifted, or if the model is biased in advance, thus cannot anticipate the performance drop in this case.
Interestingly, these problems do not occur in previous state-of-the-art, support vector machines (SVMs), which require substantially less data, enabling almost anyone to train a model from scratch. Also, as it encodes the decision boundary explicitly, SVM can reconstruct the support vectors from the training dataset using the KKT condition. Since it is a white box, one can anticipate the test’s performance in advance. In the deer classification example, if the model’s support vectors of deer have prominent antlers, using that SVM is not appropriate for photos taken in winter.
In this paper, we tackle the natural limitations of deep learning – the need for large data and black-box characteristics – by extracting SVM features in deep learning models. In doing so, we introduce the DeepKKT condition for deep models, which corresponds to the KKT condition in traditional SVMs. By either selecting deep support vectors (DSVs) from training data or generating them from already trained deep learning models, we show DSVs can play a similar role to conventional support vectors. Like support vectors can reconstruct SVM, we can reconstruct the deep models from scratch only with DSVs. Also, we show that DSVs encode the decision criterion visually, providing a global explanation for the trained deep model. Expanding beyond conventional support vectors, DSVs suggest that a trained deep classification model can also function as a latent generative model by utilizing logits as latent variables and applying DeepKKT conditions.
To this end, we generalize the KKT condition and define the DeepKKT condition considering that the data handled by a deep model is high-dimensional and multi-class. We demonstrate that the selected data points (selected DSVs) among the training data satisfying the DeepKKT condition are closer to the decision boundary than other training data, as evidenced by comparing entropy values. Also, we show that the calculated Lagrangian multiplier can reveal the level of uncertainty of the model for the sample in question. Additionally, we demonstrate that the DSVs outperform existing algorithms in the few-shot dataset distillation setting, where only a portion of the training set is used, indicating that DSVs exhibit characteristics similar to SVMs. Moreover, we confirm that modifying existing images using information obtained from DSVs allows us to change their class at will, verifying that DSVs can meaningfully explain the decision boundary. Finally, by using soft labels as latent variables in ImageNet, we generate unseen images with high fidelity.
Our contributions are as follows:
  • By generalizing the KKT condition for deep models, we propose the DeepKKT condition to extract Deep Support Vectors, which are applicable to general architectures such as ConvNet and ResNet.
  • We verify that DeepKKT can be used to extract and generate DSVs for common image datasets such as CIFAR10, CIFAR100, SVHN, and ImageNet.
  • DSVs are better than existing algorithms in few-shot dataset distillation problems.
  • DeepKKT acts as a novel model inversion algorithm that can be applied in practical situations.
  • By using the DeepKKT condition, we not only show the trained deep models can reconstruct data, but also can serve as latent generative models by using logits as latent.
@@ 2.1 SVM in Deep Learning
Numerous studies have endeavored to establish a connection between deep learning and conventional SVMs. In the theoretical side, [28] demonstrated that, in an overparameterized setting with linearly separable data, the learning dynamics of a model possessing a logistic tail are equivalent to those of a support vector machine, with the model’s normalized weights converging towards a finite value. Following this, [15] extended this equivalence to feedforward networks. This line of research relies on strong assumptions such as full-batch training and non-residual architecture without data augmentation. There exists a line of research on integrating SVM architecture into deep learning to benefit from its desirable properties. [30, 29, 27, 23, 21]. They integrated SVMs as feature extractors, extracting meaningful features in a human-crafted manner. Despite these advancements, there is still a lack of research that directly connects support vectors through this theoretical lens of equivalence. In this study, we address this theoretical equivalence of support vectors by introducing the DeepKKT condition, a KKT condition tailored for deep learning, thus applying this equivalence in a practical context. We show that reconstructing support vectors in deep models is indeed possible and obtaining high-quality support vectors is feasible under much smoother conditions.
@@ 2.2 Model Inversion Through the Lens of Maximum Margian
There is a line of research utilizing the stationarity condition, a part of the KKT condition, for model inversion. [6] firstly exploited the KKT condition for model generation, adopting SVM-like architectures. They normally conducted experiments with binary classification, a 2-layer MLP, and full-batch gradient descent. [1] extended these experiments to multi-label classification by adapting their existing architecture to a multi-class SVM structure. To ensure the generated samples lie on the data manifold, they initialized with the dataset’s mean, implying the adoption of some prior knowledge of the data. Similarly, [34] generated images through the stationarity condition, also adopting the mean-initialization and conducting experiments on the CIFAR10 [11], MNIST [4] and downsampled CelebA [14] datasets. Their work generally focused on low-dimensional, labeled datasets with a small number of classes such as CIFAR10 and MNIST, consistent with the traditional SVM setting of binary-labeled, low-dimensional datasets. In contrast, we extended our experiments to high-dimensional datasets with many classes, an area traditionally dominated by deep learning. Specifically, we conducted experiments on ImageNet [26] using a pretrained ResNet50 model following the settings described in the original paper [7].
Furthermore, previous works have concentrated on reconstructing the training dataset. In contrast, similar to generative models, our work focuses on generating unseen data from noise using a classification model. Additionally, we emphasize the original meaning of ‘support vectors’. Unlike other approaches, our Deep Support Vectors (DSVs) adhere to the traditional role of support vectors: they explain the decision criteria, and a small number of DSVs can effectively reconstruct the model.
@@ 2.3 Dataset Distillation
Dataset distillation [2, 31, 13] fundamentally aims to reduce the size of the original dataset while maintaining model performance. The achievement also involves addressing privacy concerns and alleviating communication issues between the server and client nodes. The dataset distillation problem is typically addressed under the following conditions: 1) Access to the entire dataset for gradient matching, 2) Possession of snapshots from all stages of the model’s training phase, which are impractical settings for practical usage [12]. Furthermore, these algorithms typically require Hessian computation, which imposes a heavy computational burden.
In SVM, the model can be reconstructed using support vectors. This reconstruction is more practical compared to previous dataset distillation methods, as it does not require any of the restrictive conditions. Likewise, because Deep Support Vectors (DSVs) also do not require these conditions and are Hessian-free, they can play the role of distillation under practical conditions.
@@ 3.1 Notation
In the SVM formulation, $\tilde{w}(:=[w;b])$ represents the concatenated weight vector $w$ and bias $b$. Each data instance, expanded to include the bias term, is denoted by $\tilde{x}_{i}(:=[x_{i};1])$, while the corresponding binary label is represented by $y_{i}\in\{\pm 1\}$. The Lagrange multipliers are denoted by $\alpha_{i}$’s.
Transitioning to the context of deep learning, we denote the parameter vector of a neural network model by $\theta$, which, upon optimization, yields $\theta^{*}$ as the set of learned weights. The mapping function $\Phi(x_{i};\theta)$ represents the transformation of input data into a $C$-dimensional logit in a $C$-class classification problem in a manner dictated by the parameters $\theta$, i.e., $\Phi(x_{i};\theta)=[\Phi_{1}(x_{i};\theta),\cdots,\Phi_{C}(x_{i};\theta)]^{T}\in\mathbb{R}^{C}$. We define the score as the logit of a target class, i.e., $\Phi_{y_{i}}(x_{i};\theta)$. If the score is the largest among logits, i.e., $\operatorname*{arg\,max}_{c}\ \Phi_{c}(x_{i};\theta)=y_{i}$, then it correctly classifies the sample. The Lagrange multipliers adapted to the optimization in deep learning are represented by $\lambda_{i}$’s. $(x,y)$ denotes a pair of input and output and $\mathcal{I}$ is the index set with $|\mathcal{I}|=n$.
@@ 3.2 Support Vector Machines
The fundamental concept of Support Vector Machines (SVMs) is to find the optimal hyperplane that classifies the given data. This hyperplane is defined by the closest data points to itself known as support vectors, and the distance between the support vectors and the hyperplane is termed the margin. The hyperplane must classify the classes correctly while maximizing the margin. This leads to the following KKT conditions that an SVM must satisfy: (1) Primal feasibility: $\forall i,\ y_{i}\tilde{w}^{T}\tilde{x}_{i}\geq 1$, (2) Dual feasibility: $\forall i,\ \alpha_{i}\geq 0$, (3) Complementary slackness: $\alpha_{i}\left(y_{i}\tilde{w}^{T}\tilde{x}_{i}-1\right)=0$ and (4) Stationarity: $\tilde{w}=\sum_{i=1}^{n}\alpha_{i}y_{i}\tilde{x}_{i}$.
The primal and dual conditions ensure these critical values correctly classify the data while being outside the margin. The complementary slackness condition mandates that support vectors lie on the decision boundary. The stationarity condition ensures that the derivative of the Lagrangian is zero.
The final condition, stationarity, offers profound insights into SVMs. It underscores that support vectors encode the decision boundary $\tilde{w}$. Consequently, identifying the decision hyperplane in SVMs is tantamount to pinpointing the corresponding support vectors. This implies that with a trained model at our disposal, we can reconstruct the SVM in two distinct ways:
1. Support Vector Selection: From the trained model, we can extract support vectors among the training data that inherently encode the decision hyperplane.
@@ 4 Deep Support Vector
This section presents the specific conditions that DSVs (Deep Support Vectors) must satisfy and discusses how to get an optimization loss to meet these conditions.
@@ 4.1 DeepKKT
$\begin{split}&\text{Hinge Loss:}\ L_{h}(x_{i},y_{i};\tilde{w})=\frac{1}{n}\sum_{i=1}^{n}\max(0,1-y_{i}(\tilde{w}^{T}\tilde{x}_{i})),\\ &\quad\nabla_{\tilde{w}}L_{h}=\frac{1}{n}\sum_{i=1}^{n}\begin{cases}0,&\text{if }y_{i}(\tilde{w}^{T}\tilde{x}_{i})\geq 1\\ -y_{i}\tilde{x}_{i},&\text{otherwise.}\end{cases}\end{split}$ (1)
With Eq. (LABEL:eq:hingegrad), the stationarity condition of SVM becomes
$w^{*}:=-\sum_{i=1}^{n}\alpha_{i}\nabla_{w}L_{{\color[rgb]{0,0,0}h}}(x_{i},y_{i};w^{*}),\quad\text{s.t.}\ \alpha_{i}\geq 0.$ (2)
@@
In this paper, we extend the KKT conditions to deep learning. In doing so, the two main hurdles of a deep network different from a linear binary SVM are 1) the nonlinearity of $\Phi(x_{i};\theta)$ taking the role of $\tilde{w}^{T}\tilde{x}_{i}$ and 2) multi-class nature of a deep learning model.
Considering that the role of the primal feasibility condition is to correctly classify $x_{i}$ into $y_{i}$, we can enforce the score $\Phi_{y_{i}}(x_{i};\theta)$ for the correct class $y_{i}$ to take the maximum value among all the logits with some margin $\epsilon$, i.e.,
@@
We can relax this discontinuity with a continuous surrogate function $-L(\Phi(x_{i};\theta^{*}),y_{i})$, which is the negative loss function to maximize. Note that if we take the cross-entropy loss for $L$, it becomes
@@
$\theta^{*}=-\sum_{i=1}^{n}\lambda_{i}\nabla_{\theta^{*}}L(\Phi(x_{i};\theta^{*}),y_{i}).$ (5)
This condition is a generalized formulation of Eq. (2), where we substitute the linear model $\tilde{w}^{T}\tilde{x}$ with a nonlinear model $\Phi(x;\theta)$, and the binary classification hinge loss $L_{h}$ with multi-class classification loss $L$. Furthermore, the stationarity condition reflects the dynamics of overparameterized deep learning models. We provide an analogy with respect to [28] in Appendix C.
However, these conditions are not enough for deep learning. As mentioned before, we are interested in dealing with high dimensional manifolds. Compared to the problems dealt with in the classical SVMs, the input dimensions of deep learning problems are typically much higher. In this case, the data are likely to lie along a low-dimensional latent manifold $\mathcal{M}$ inside the high-dimensional space. To make a generated sample be a plausible DSV, it not only should satisfy the generalized KKT condition but also should lie in an appropriate data manifold, i.e., $x\in\mathcal{M}$.
@@
Primal feasibility: $\forall\,i\in\mathcal{I},\quad\operatorname*{arg\,max}_{c}\;\Phi_{c}(x_{i};\theta^{*})=y_{i}$ (6)
@@
4.2 Deep Support Vectors, From Dust to Diamonds
Sec. 4.1 explored the KKT conditions in the context of deep learning and how these conditions can be used to formulate a loss function. It is important to note that our goal is not to construct the model $\Phi$, but rather to generate support vectors of an already-trained model $\Phi$ with its parameter $\theta^{*}$.
To reconstruct support vectors from a trained deep learning model, sampling or synthesizing support vectors is essential. By replacing the optimization variable $\theta$ with $x$, we shift our focus to utilizing the DeepKKT conditions for generating or evaluating input $x$ rather than $\theta$. This adjustment necessitates a consideration of the data’s inherent characteristics, specifically its multiclass nature and the tendency of the data to reside on a lower-dimensional manifold in an ambient space.